### v1

# Counting integer points on affine surfaces with a side condition

Abstract.
We extend work of Heath-Brown and Salberger, based on the determinant method, to provide a uniform upper bound for the number of integral points of bounded height on an affine surface, which are subject to a polynomial congruence condition. This is applied to get a new uniform bound for points on diagonal quadric surfaces, and to a problem about the representation of integers as a sum of four unlike powers.
2020 Mathematics Subject Classification:
11D45 (11D09, 11D41, 11P05)
1. Introduction
Let $f\in\mathbb{Z}[x_{1},x_{2},x_{3}]$ be an absolutely irreducible polynomial of degree $d$. The determinant method of Bombieri–Pila [1] has proved a tremendously useful tool for bounding the number $N(f;B)$ of zeros $\underline{x}=(x_{1},x_{2},x_{3})\in\mathbb{Z}^{3}$ of $f$ with $|x_{1}|,|x_{2}|,|x_{3}|\leqslant B$, for a large parameter $B$. As noticed by Pila [12], it is possible to apply the bounds in [1] to prove that $N(f;B)=O_{d,\varepsilon}(B^{1+1/d+\varepsilon})$ for any $\varepsilon>0$, where the implied constant only depends on $d$ and $\varepsilon$. The determinant method has been extensively developed, most notably by Heath-Brown [7, 8] and by Salberger [14]. It follows from Salberger’s work [14, Cor. 7.3] that
$N(f;B)=O_{d,\varepsilon}(B^{2/\sqrt{d}+\varepsilon}+B^{1+\varepsilon}),$ (1.1)
which is now the state of the art for this problem.
In this paper we shall study a variant of this counting problem in which lopsided boxes are allowed, and moreover, in which an additional side condition is imposed, in the form of a congruence. Let $g\in\mathbb{Z}[x_{1},x_{2},x_{3}]$ be a further non-constant polynomial and let $q\in\mathbb{N}$. Under suitable hypotheses we shall investigate the size of the set
$S_{q}(f,g;\underline{B})=\left\{\underline{x}\in\mathbb{Z}^{3}:|x_{i}|\leqslant B_{i},~{}f(\underline{x})=0,~{}g(\underline{x})\equiv 0\bmod{q}\right\},$
where $\underline{B}=(B_{1},B_{2},B_{3})$, with $B_{1},B_{2},B_{3}\geqslant 2$. We simply write $S_{q}(f,g;B)$, when $B_{1}=B_{2}=B_{3}=B$.
We will proceed under the assumption that $g$ does not depend on $x_{1}$, so that $g\in\mathbb{Z}[x_{2},x_{3}]$. When $g$ is homogeneous, one can cover the congruence condition $g(\underline{x})\equiv 0\bmod{q}$ by a small number of lattices. Techniques from the geometry of numbers then reduce the problem to bounding #$\#S_{1}(\tilde{f},0;\underline{C})$ for a suitable degree $d$ polynomial $\tilde{f}$ and a new triple $\underline{C}=(C_{1},C_{2},C_{3})$ that depends on $q$. The main aim of this paper is to harness arguments of Heath-Brown [7, 8] and Salberger [14] to extract information about the case that $g$ is not necessarily homogeneous. We shall apply our work to get new uniform bounds for the density of integral points on the quadric surface $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}=n$, and on the threefold $x_{1}^{k}+x_{2}^{l}+x_{3}^{m}+x_{4}^{k}=n$.
For given $f\in\mathbb{Z}[x_{1},x_{2},x_{3}]$ and $g\in\mathbb{Z}[x_{2},x_{3}]$ we will need to introduce some notation and conventions that will stay in place throughout this paper. We will allow all implied constants to depend on the degrees of $f$ and $g$. Let $\underline{B}\in\mathbb{R}^{3}$ with $B_{1},B_{2},B_{3}\geqslant 2$. We fix a strict total order $\prec$ on $\mathbb{Z}_{\geqslant 0}^{3}$. We shall assume that it is linear, so that $\underline{a}\prec\underline{b}$ and $\underline{c}\prec\underline{d}$ implies that $\underline{a}+\underline{c}\prec\underline{b}+\underline{d}$, for any $\underline{a},\underline{b},\underline{c},\underline{d}\in\mathbb{Z}_{\geqslant 0}^{3}$. We call any such order a linear total order. The lexicographic order is an example of a linear total order (and it is this order that we shall use in our applications). Among the monomials $\underline{x}^{\underline{m}}=x_{1}^{m_{1}}x_{2}^{m_{2}}x_{3}^{m_{3}}$ that appear in $f$ with a non-zero coefficient, we choose $\underline{m}\in\mathbb{Z}_{\geqslant 0}^{3}$ to be the maximum for the order. We are going to work under the assumption that $\underline{m}\neq\underline{0}$. For this choice of $\underline{m}$, we put
$T_{\underline{m}}=\underline{B}^{\underline{m}}.$
Let
$S=\max_{\underline{e}}\{\log\underline{B}^{\underline{e}}\},$
where the maximum runs over exponents $\underline{e}\in\mathbb{Z}_{\geqslant 0}^{3}$ such that $\underline{x}^{\underline{e}}$ appears in $g$ with non-zero coefficient. (Note that $e_{1}=0$ in any such exponent vector by our assumption on $g$.) Let
$B^{\prime}=\min\{B_{1},B_{2},B_{3}\},\quad B=\max\{B_{1},B_{2},B_{3}\},$ (1.2)
and
$K=\exp\left(\sqrt{\frac{\log B_{1}\log B_{2}\log B_{3}}{\log T_{\underline{m}}}}\left(1-\frac{m_{1}\log B_{1}}{2S\log T_{\underline{m}}}\log q\right)\right).$ (1.3)
Note that $\log T_{\underline{m}}$ and $S$ are not zero (since $\underline{m}\neq\underline{0}$ and $g$ is non-constant) and then $K$ is well-defined. Then, with this notation, the following will be proved in Section 2.
Theorem 1.1.
Let $\varepsilon>0$ and $q\in\mathbb{N}$. Let $f\in\mathbb{Z}[x_{1},x_{2},x_{3}]$ be absolutely irreducible and let $g\in\mathbb{Z}[x_{2},x_{3}]$. Assume that $\gcd(q,g(0,0))=1$ and let $B,B^{\prime},K$ be given by (1.2) and (1.3). Then there exist $f_{1},\ldots,f_{J}\in\mathbb{Z}[x_{1},x_{2},x_{3}],$ and a finite collection of points $Z\subset V(f)$, such that the following hold:
  1. (1)
    $J=O_{\varepsilon}(\max\{1,KB^{\varepsilon}\})$;
  2. (2)
    each $f_{j}$ is coprime to $f$ and has degree $O_{\varepsilon}(\log B/\log B^{\prime})$, for $j\leqslant J$;
  3. (3)
    #$\#Z=O_{\varepsilon}(\max\{1,KB^{\varepsilon}\}^{2})$; and
  4. (4)
    for each $\underline{x}\in S_{q}(f,g;\underline{B})\setminus Z$, there exists $j\leqslant J$ such that
    $f(x_{1},x_{2},x_{3})=f_{j}(x_{1},x_{2},x_{3})=0.$
We emphasise that the implied constants in this result only depend on the choice of $\varepsilon$ and on the degrees of $f$ and $g$, a convention that will remain in place throughout this paper. The proof relies crucially on the breakthrough work of Salberger [14]. In order to find the best possible bound, one chooses the linear total order that minimises $K$. When $q=1$, we may assume without loss of generality that the values of $\log B_{i}$ are linearly independent over $\mathbb{Q}$. One may then check that the order $\underline{a}\prec\underline{b}$ if and only if $\underline{B}^{\underline{a}}<\underline{B}^{\underline{b}}$ is a linear total order on $\mathbb{Z}_{\geqslant 0}^{3}$. Thus we may take
$K=\exp\left(\sqrt{\frac{\log B_{1}\log B_{2}\log B_{3}}{\max_{\underline{m}}T_{\underline{m}}}}\right)$
in (1.3), where the maximum is taken over all exponents $\underline{m}\in\mathbb{Z}_{\geqslant 0}^{3}$ such that $\underline{x}^{\underline{m}}$ appears in $f$ with non-zero coefficient. Thus Theorem 1.1 recovers work of the first author [2, Lemma 1], in which a similar result without the side condition was already extracted from [8] and [14]. (In fact, Theorem 1.1 corrects a mistake in [2, Lemma 1], in which it is claimed that $f_{j}$ has degree $O_{\varepsilon}(1)$ for all $j\leqslant J$.) This in turn could be used to recover the bound (1.1) of Salberger when $B_{1}=B_{2}=B_{3}=B$.
In the case $B_{1}=B_{2}=B_{3}=B$ we can relax the condition $\gcd(q,g(0,0))=1$ somewhat. In this case (1.3) becomes
$K=B^{\frac{1}{\sqrt{m_{1}+m_{2}+m_{3}}}}q^{-\frac{m_{1}}{2(m_{1}+m_{2}+m_{3})^{3/2}\deg g}},$ (1.4)
for any $\underline{m}\in\mathbb{Z}_{\geqslant 0}^{3}$ that is a maximum (among the exponents of $f$) for a fixed linear total order. We shall also prove the following result in Section 2.
Theorem 1.2.
Let $\varepsilon>0$ and $q\in\mathbb{N}$. Let $f\in\mathbb{Z}[x_{1},x_{2},x_{3}]$ be absolutely irreducible and let $g\in\mathbb{Z}[x_{2},x_{3}]$. Let $K$ be given by (1.4) and assume that
$\gcd(q,g(0,0),g_{0}(1,0),g_{0}(0,1))=1,$ (1.5)
where $g_{0}$ is the top degree part of $g$. Then there exist $f_{1},\ldots,f_{J}\in\mathbb{Z}[x_{1},x_{2},x_{3}],$ and a finite collection of points $Z\subset V(f)$, such that the following hold:
  1. (1)
    $J=O_{\varepsilon}(\max\{1,KB^{\varepsilon}\})$;
  2. (2)
    each $f_{j}$ is coprime to $f$ and has degree $O_{\varepsilon}(1)$, for $j\leqslant J$;
  3. (3)
    #$\#Z=O_{\varepsilon}(\max\{1,KB^{\varepsilon}\}^{2})$; and
  4. (4)
    for each $\underline{x}\in S_{q}(f,g;B)\setminus Z$, there exists $j\leqslant J$ such that
    $f(x_{1},x_{2},x_{3})=f_{j}(x_{1},x_{2},x_{3})=0.$
We now collect two applications of our bound for surfaces with a side condition. The first concerns a uniform upper bound for the affine quadric surface $S\subset\mathbb{A}^{3}$ defined by $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}=n$, for suitable coefficients $a_{1},a_{2},a_{3},n\in\mathbb{Z}$. Let
$N_{1}(B)=\#\{\underline{x}\in S(\mathbb{Z}):|\underline{x}|\leqslant B\},$ (1.6)
where $|\underline{x}|=\max\{|x_{1}|,|x_{2}|,|x_{3}|\}$ is the sup-norm. In some applications it is useful to have a bound for $N_{1}(B)$ that gets better when $|\underline{a}|$ gets larger. When $n=0$ and the integer solutions are further restricted to be primitive, such a bound is provided in [4, Cor. 2] using arguments from the geometry of numbers. We are interested here in the case $n\neq 0$. In this case a divisor function bound (in the form [3, Thm. 1.11], for example) readily yields
$N_{1}(B)\ll_{\varepsilon}B^{1+\varepsilon},$
for any $\varepsilon>0$, where the implied constant only depends on $\varepsilon>0$. This bound is optimal up to the factor $B^{\varepsilon}$, since $S$ can contain rational lines if $-a_{1}a_{2}a_{3}n=\square$. We are interested in obtaining bounds for $N_{1}(B)$ that improve on this when $|\underline{a}|$ is large and $-a_{1}a_{2}a_{3}n\neq\square$. We shall build on the proof of Theorem 1.2 to prove the following estimate in Section 3, which improves on the divisor bound as soon as $|\underline{a}|>B^{1+\delta}$, for any fixed $\delta>0$.
Corollary 1.3.
Let $a_{1},a_{2},a_{3},n\in\mathbb{Z}$, with $-a_{1}a_{2}a_{3}n\neq\square$ and $\gcd(a_{1},a_{2},a_{3},n)=1$. Then, for all $\varepsilon>0$, we have
$N_{1}(B)\ll_{\varepsilon}\frac{B^{\frac{4}{3}+\varepsilon}}{|\underline{a}|^{1/3}}+\frac{B^{\frac{7}{6}+\varepsilon}}{|\underline{a}|^{1/6}}+B^{\frac{1}{2}+\varepsilon}.$
A direct application of Theorem 1.2 would actually lead to the weaker bound $N_{1}(B)\ll_{\varepsilon}|\underline{a}|^{-1/(2\sqrt{2})}B^{\sqrt{2}+\varepsilon}+|\underline{a}|^{-1/(4\sqrt{2})}B^{\frac{1}{\sqrt{2}}+\frac{1}{2}+\varepsilon}+B^{\frac{1}{2}+\varepsilon}.$ In order to arrive at the statement of Corollary 1.3, we shall exploit earlier work of Salberger [13] that is more efficient for surfaces of degree $2$ or $3$. Note that a bound for $N_{1}(B)$ can also be extracted from work of Ellenberg and Venkatesh [6, Prop. 1 and Rem. 2], which uses the determinant method to get better upper bounds when the underlying polynomial has large coefficients. When specialised to our setting (with $d=2$ and $n=2$), their work shows that the set of $\underline{x}\in S(\mathbb{Z})$ with $|\underline{x}|\leqslant B$ are covered by $O_{\varepsilon}(|\underline{a}|^{-1/(2\sqrt{2})}B^{\sqrt{2}+\varepsilon}+B^{\varepsilon})$ curves of degree $O_{\varepsilon}(1)$. This was proved before Salberger’s work [13, 14] emerged and leads to a strictly weaker bound for $N_{1}(B)$.
Given non-zero $N\in\mathbb{Z}$, our next application is to the problem of estimating
$N_{2}(B)=\#\{\underline{x}\in\mathbb{Z}^{4}:|\underline{x}|\leqslant B,~{}x_{1}^{k}+x_{2}^{l}+x_{3}^{m}+x_{4}^{k}=N\},$
for suitable integer exponents $k,l,m\geqslant 2$. Combining divisor bounds with uniform bounds for Thue equations, one can easily show that $N_{2}(B)=O_{\varepsilon}(B^{2+\varepsilon}|N|^{\varepsilon})$. The problem of improving this bound has a rich history, but most work focuses on counting representations of a large positive integer $N$ by $\underline{x}\in\mathbb{Z}_{\geqslant 0}^{4}$. Let $N_{2}^{+}(B)$ denote this counting function. Then an old (and still unbeaten) result of Hooley [9] provides the upper bound $N_{2}^{+}(B)=O_{\varepsilon}(B^{2-1/6+\varepsilon})$ in the case $k=l=m=3$. Hooley used sieve methods, an approach that Wisdom extended to handle the case $k=l=m=5$ in [15], and the cases $k=l=3$ and $m\in\{4,5\}$ in [16]. More recently, Salberger’s version of the determinant method has also been applied to this problem by Marmon [10], with the outcome that $N_{2}^{+}(B)\ll_{\varepsilon}B^{1+2/\sqrt{k}+\varepsilon},$ when $k=l=m$. (Marmon gets better bounds when the implied constant is allowed to depend additionally on $N$, but our focus in this paper is on uniform bounds.) We shall deduce the following result from Theorem 1.2 in Section 4.
Corollary 1.4.
Let $N\in\mathbb{Z}_{\neq 0}$ and let $k,l,m$ be integers such that $k\geqslant 13$ is odd and $k>l>m\geqslant 2$. Then, for all $\varepsilon>0$, we have
$N_{2}(B)\ll_{\varepsilon}B^{\frac{4}{3}+\frac{1}{\sqrt{k-1}}(1-\frac{1}{2l})+\varepsilon}+B^{1+\frac{2}{\sqrt{k-1}}(1-\frac{1}{2l})+\varepsilon}.$
We shall see that the restriction $k\geqslant 13$ allows us to control the contribution from lines and conics in the hypersurface through work of Newman and Slater [11]. The assumption that $k$ is odd allows us to factorise $x_{1}^{k}+x_{4}^{k}$ over $\mathbb{Q}$, which is a crucial first step in the proof.
The determinant method approach of Marmon [10] could also be used to give an upper bound for $N_{2}(B)$. His approach is based on fixing a variable $x_{3}$, say, and then applying the argument behind (1.1) to estimate the number of points that contribute to $N_{2}(B)$ on the resulting affine surface. When carried through this would lead to an alternative bound
$N_{2}(B)\ll_{\varepsilon}B^{\frac{4}{3}+\frac{1}{\sqrt{k}}+\varepsilon}+B^{1+\frac{2}{\sqrt{k}}+\varepsilon},$
for $k\geqslant l\geqslant m\geqslant 2$, with $k\geqslant 13$. While this is more efficient for $k=l$, Corollary 1.4 is stronger for $k>l$, since then
$\frac{1}{\sqrt{k-1}}\left(1-\frac{1}{2l}\right)<\frac{1}{\sqrt{k}}.$
Acknowledgements
The authors are grateful to Jakob Glas for helpful conversations and to Per Salberger for comments that led to an improvement in Corollary 1.3. While working on this paper the first author was supported by a FWF grant (DOI 10.55776/P36278) and the second author was supported by the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie Grant Agreement No. 101034413.
2. Salberger’s global determinant method with a side condition
We begin by recalling the notation that was introduced in the introduction. Let $f\in\mathbb{Z}[x_{1},x_{2},x_{3}]$ be absolutely irreducible, let $g\in\mathbb{Z}[x_{2},x_{3}]$ and let $B_{1},B_{2},B_{3}\geqslant 2$. Let $\underline{m}\in\mathbb{Z}_{\geqslant 0}^{3}\setminus\{\underline{0}\}$ be the maximum among exponents appearing in $f$ for a fixed linear total order. Then we put $T_{\underline{m}}=\underline{B}^{\underline{m}}$ and we define $S=\max_{\underline{e}}\{\log\underline{B}^{\underline{e}}\}$, where the maximum runs over exponents $\underline{e}\in\mathbb{Z}_{\geqslant 0}^{3}$ such that $\underline{x}^{\underline{e}}$ appears in $g$ with a non-zero coefficient. Let $B$ and $B^{\prime}$ be given by (1.2). Let $\varepsilon>0$. It will be convenient to define
$K_{\varepsilon}=\exp\left(\sqrt{\frac{\prod_{i}\log B_{i}}{\log T_{\underline{m}}}}\left(1-\frac{m_{1}\log B_{1}}{2S\log T_{\underline{m}}}\log q\right)+\varepsilon\log B\right),$ (2.1)
so that $K_{\varepsilon}=KB^{\varepsilon}$ in the notation of (1.3).
Our proof of Theorems 1.1 and 1.2 will be achieved by blending arguments of Heath-Brown [8] with the relevant arguments of Salberger [14]. A singular point of $f(\underline{x})=0$ satisfies $\frac{\partial f}{\partial x_{i}}(\underline{x})=0$ for $1\leqslant i\leqslant 3$. Not all of these derivatives can vanish identically on the zero locus of $f$, since $f$ is absolutely irreducible. Thus we can account for all singular points if we include a non-zero partial derivative among the auxiliary polynomials $f_{1},\dots,f_{J}$.
We proceed by introducing the exponent set that will play a critical role in our argument. For $Y\geqslant 2$, let
$\mathcal{E}(Y)=\{\underline{e}\in\mathbb{Z}_{\geqslant 0}^{3}:\log\underline{B}^{\underline{e}}\leqslant Y,e_{i}<m_{i}\text{ for at least one }i\}$
and put
$\mathcal{E}^{1}(Y)=\{\underline{e}\in\mathcal{E}(Y):e_{1}<m_{1}\}.$ (2.2)
Let $h(\underline{x})=\sum_{\underline{e}\in\mathcal{E}(Y)}a_{\underline{e}}\underline{x}^{\underline{e}}$, for any coefficients $a_{\underline{e}}\in\mathbb{Z}$. We proceed to show that $h$ must be coprime to $f$. Indeed, suppose for a contradiction that $h=fk$, for some further polynomial $k\in\overline{\mathbb{Q}}[\underline{x}]$. Let $\underline{k}$ be the maximum exponent (for the fixed linear total order) that appears in $k$ with non-zero coefficient. Note that if $\underline{m^{\prime}}\prec\underline{m}$ and $\underline{k^{\prime}}\prec\underline{k}$ then $\underline{m^{\prime}}+\underline{k^{\prime}}\prec\underline{m}+\underline{k}$. Hence it follows that the exponent $\underline{m}+\underline{k}$ can be written in a unique way as the sum of an exponent in $f$ and an exponent in $k$, since $\underline{m}$ and $\underline{k}$ are the maxima for the linear total order. Thus $\underline{m}+\underline{k}$ must appear as an exponent in $h$ with a non-zero coefficient, whence $\underline{m}+\underline{k}\in\mathcal{E}(Y)$. But $m_{i}+k_{i}\geqslant m_{i}$ for each $1\leqslant i\leqslant 3$, which contradicts the definition of $\mathcal{E}(Y)$. Thus we cannot have that $h$ is a multiple of $f$ over $\overline{\mathbb{Q}}$.
We let $E=\#\mathcal{E}(Y)$. For each $Z\geqslant\log T_{\underline{m}}$, it follows from [8, Eqs. (5.9) and (5.10)] that there exists $Y\in[Z,2Z]$ such that
$E=\frac{\log T_{\underline{m}}}{\prod_{i}\log B_{i}}\frac{Y^{2}}{2}\left(1+O\left(\sqrt{\frac{\log T_{\underline{m}}}{Y}}\right)\right)$ (2.3)
and
$\sum_{\underline{e}\in\mathcal{E}(Y)}\log\underline{B}^{\underline{e}}=\frac{\log T_{\underline{m}}}{\prod_{i}\log B_{i}}\frac{Y^{3}}{3}\left(1+O\left(\sqrt{\frac{\log T_{\underline{m}}}{Y}}\right)\right).$ (2.4)
We fix such a $Y$ with $Y\gg\log B$, for a sufficiently large implied constant that depends only on the degrees of $f$ and $g$. In particular, we must then have $Y\gg\log T_{\underline{m}}$, since $T_{\underline{m}}\leqslant B^{\deg f}$.
Let $q\in\mathbb{N}$. We henceforth write $X\subset\mathbb{A}^{3}$ for the geometrically irreducible surface defined by $f=0$. We redefine $S_{q}(f,g;\underline{B})$ to focus on non-singular points of $X$, so that $\nabla f(\underline{x})\neq\underline{0}$ for any $\underline{x}\in S_{q}(f,g;\underline{B})$. Let $r_{1}<\dots<r_{u}$ be a sequence of primes that are coprime with $q$ and such that the reduction modulo $r_{i}$ of $X$ is geometrically irreducible. Let
$r=\begin{cases}r_{1}\cdots r_{u}&\text{if $u\geqslant 1$,}\\ 1&\text{if $u=0$.}\end{cases}$ (2.5)
For each $i\leqslant u$, let $\underline{t}_{i}\in X(\mathbb{F}_{r_{i}})$ be a non-singular point of the reduction of $X$ modulo $r_{i}$. Let $\underline{t}=(\underline{t}_{1},\dots,\underline{t}_{u})$ and put
$S_{r,\underline{t}}=\left\{\underline{x}\in X(\mathbb{Z}):\begin{array}[]{l}\underline{x}\equiv\underline{t}_{i}\bmod{r_{i}}~{}\forall i\leqslant u\\ |x_{1}|\leqslant B_{1},|x_{2}|\leqslant B_{2},|x_{3}|\leqslant B_{3}\\ g(\underline{x})\equiv 0\bmod q\end{array}\right\}.$ (2.6)
Let $J=\#S_{r,\underline{t}}$ and let $\underline{x}_{1},\dots,\underline{x}_{J}$ be the elements of $S_{r,\underline{t}}$. We may then form the $J\times E$ matrix
$\mathbf{M}=\left(\underline{x}_{j}^{\underline{e}}\right)_{\begin{subarray}{c}\underline{e}\in\mathcal{E}(Y)\\ 1\leqslant j\leqslant J\end{subarray}},$ (2.7)
with the goal being to prove that $\operatorname{rank}\mathbf{M}<E$ under suitable choices for $r$ and $Y$. To achieve this, we may clearly proceed under the assumption that $J\geqslant E$ and, on relabelling, we may work with the principal $E\times E$ minor
$\Delta=\left\lvert\det\left(\underline{x}_{j}^{\underline{e}}\right)_{\begin{subarray}{c}\underline{e}\in\mathcal{E}(Y)\\ 1\leqslant j\leqslant E\end{subarray}}\right\rvert.$
Our goal is now to show that $\Delta=0$, which we shall achieve by comparing upper bounds for $\Delta$ with information about the $p$-adic valuation of $\Delta$ for many primes $p$.
Note that $|\underline{x}_{j}^{\underline{e}}|\leqslant B_{1}^{e_{1}}B_{2}^{e_{2}}B_{3}^{e_{3}}$, for each $1\leqslant j\leqslant J$. Hence it follows from (2.3) and (2.4) that
$\log\Delta\leqslant\log E^{E}+\sum_{\underline{e}\in\mathcal{E}(Y)}\log\underline{B}^{\underline{e}}=\frac{\log T_{\underline{m}}}{\prod_{i}\log B_{i}}\frac{Y^{3}}{3}\left(1+O\left(\sqrt{\frac{\log T_{\underline{m}}}{Y}}\right)\right).$ (2.8)
Before turning to the $p$-adic valuation of $\Delta$, it will be convenient to record the estimate
$Y^{3}=2\sqrt{2}E^{3/2}\left(\frac{\prod_{i}\log B_{i}}{\log T_{\underline{m}}}\right)^{3/2}\left(1+O\left(\sqrt{\frac{\log T_{\underline{m}}}{Y}}\right)\right),$ (2.9)
which easily follows from (2.3) and the assumption that $Y\gg\log T_{\underline{m}}$, for a sufficiently large implied constant.
We are now ready to collect together estimates for the divisibility of $\Delta$ by various prime powers. Let $r_{1}<\cdots<r_{u}$ be the primes that were chosen above and recall the definition (2.5) of $r$. The following result is due to Heath-Brown [8, Eq. (5.11)], combined with the estimate in (2.9) for $Y^{3}$.
Lemma 2.1.
There exists $\nu\in\mathbb{N}$ such that $r^{\nu}\mid\Delta$, with
$\nu=\frac{2\sqrt{2}E^{\frac{3}{2}}}{3}\left(1+O\left(\sqrt{\frac{\log T_{\underline{m}}}{Y}}\right)\right).$
The following result is due to Salberger [14, Lemmas 1.4 and 1.5].
Lemma 2.2.
Let $p$ be a prime and assume that the reduction modulo $p$ of $X$ is geometrically integral. There exists $N_{p}\in\mathbb{N}$ such that $p^{N_{p}}\mid\Delta$, with
$N_{p}\geqslant\frac{2\sqrt{2}}{3}\frac{E^{3/2}}{p}+O\left(E+E^{3/2}p^{-3/2}\right).$
We shall need a good lower bound for the average of $N_{p}$, weighted by $\log p$.
Lemma 2.3.
Let $\pi_{X}$ be the product of primes $p$ such that the reduction modulo $p$ of $X$ is not geometrically integral. Let $N_{p}$ be as in Lemma 2.2, for any prime $p$. Then
$\sum_{\begin{subarray}{c}p\leqslant x\\ p\nmid qr\pi_{X}\end{subarray}}N_{p}\log p\geqslant\frac{2\sqrt{2}}{3}E^{3/2}\left(\log x+O\left(\frac{x}{\sqrt{E}}+\log\log 3qr\pi_{X}\right)\right).$
Proof.
It follows from Lemma 2.2 that
$\sum_{\begin{subarray}{c}p\leqslant x\\ p\nmid qr\pi_{X}\end{subarray}}N_{p}\log p$ $\geqslant\sum_{\begin{subarray}{c}p\leqslant x\\ p\nmid qr\pi_{X}\end{subarray}}\left(\frac{2\sqrt{2}}{3}\frac{E^{3/2}\log p}{p}+O\left(E\log p+\frac{E^{3/2}\log p}{p^{3/2}}\right)\right)$
$=\frac{2\sqrt{2}}{3}E^{3/2}\sum_{\begin{subarray}{c}p\leqslant x\\ p\nmid qr\pi_{X}\end{subarray}}\frac{\log p}{p}+O\left(Ex+E^{3/2}\right),$
by the prime number theorem. According to [14, Lemma 1.10], we have
$\sum_{\begin{subarray}{c}p\leqslant x\\ p\nmid qr\pi_{X}\end{subarray}}\frac{\log p}{p}=\sum_{\begin{subarray}{c}p\leqslant x\end{subarray}}\frac{\log p}{p}+O(\log\log 3qr\pi_{X})=\log x+O(\log\log 3qr\pi_{X}).$
Hence the statement of the lemma follows. ∎
Our remaining task is to show that $q^{\lambda}$ divides $\Delta$ for suitable $\lambda\in\mathbb{N}$, using the congruence condition in $S_{q}(f,g;\underline{B})$. This is the part of our paper that relies crucially on the fact that $g$ does not depend on $x_{1}$. Indeed, our approach will use the fact that all the monomials appearing in $g(x_{2},x_{3})^{m}$ belong to $\mathcal{E}(Y)$, provided that $m$ is sufficiently small, a property that could fail if $g$ were to depend on $x_{1}$, since it could obstruct the constraint that $e_{i}<m_{i}$ for at least one index $i\in\{1,2,3\}$.
In what follows, it will be convenient to set
$R=\frac{m_{1}(\log B_{1})^{\frac{3}{2}}(\log B_{2})^{\frac{1}{2}}(\log B_{3})^{\frac{1}{2}}}{2S(\log T_{\underline{m}})^{\frac{3}{2}}},$ (2.10)
where $T_{\underline{m}}$ and $S$ were defined at the start of the section. In the light of (2.1), we clearly have
$\log K_{\varepsilon}=\sqrt{\frac{\prod_{i}\log B_{i}}{\log T_{\underline{m}}}}-R\log q+\varepsilon\log B.$ (2.11)
Recall the definition (2.2) of the set $\mathcal{E}^{1}(Y)$. For any $\underline{e}\in\mathcal{E}^{1}(Y)$ and $\underline{t}\in\mathbb{Z}_{\geqslant 0}^{3}$, let $\lambda_{\underline{e},\underline{t}}$ be the largest integer $\lambda\in\mathbb{Z}_{\geqslant 0}$ such that $\underline{e}$ can be written as $\underline{e}=\underline{e}^{\prime}+\lambda\underline{t}$, for some $\underline{e}^{\prime}\in\mathcal{E}^{1}(Y)$. Note that $0\leqslant\lambda_{\underline{e},\underline{t}}\leqslant\max\{e_{1},e_{2},e_{3}\}$ if $\underline{t}\neq\underline{0}$. If $\underline{t}=\underline{0}$, then we put $\lambda_{\underline{e},\underline{t}}=\infty$. Bearing this notation in mind, we proceed to prove the following result.
Lemma 2.4.
Suppose that $q$ is a prime power and let $\underline{t}=(0,t_{2},t_{3})$ be such that the coefficient of $\underline{x}^{\underline{t}}$ in $g$ is non-zero and coprime to $q$. Then $q^{\lambda}\mid\Delta$, where
$\lambda=\sum_{\underline{e}\in\mathcal{E}^{1}(Y)}\min\left\{\lambda_{\underline{e},\underline{t}},\left\lfloor\frac{Y-\log\underline{B}^{\underline{e}}}{S-\log\underline{B}^{\underline{t}}}\right\rfloor\right\}.$ (2.12)
Proof.
Let $c_{\underline{t}}$ denote the coefficient of $\underline{x}^{\underline{t}}$ in $g$, which is assumed to be coprime to $q$. Let $z\in\mathbb{Z}$ such that $zc_{\underline{t}}\equiv 1\bmod q$ and write $s=(zc_{\underline{t}}-1)/q\in\mathbb{Z}$. Further, let $g_{q}(\underline{x})=zg(\underline{x})-qs\underline{x}^{\underline{t}}$. Then $g_{q}\in\mathbb{Z}[x_{2},x_{3}]$ and the coefficient of $\underline{x}^{\underline{t}}$ is $1$. It is clear that
$g_{q}(\underline{x})\equiv 0\bmod q$ (2.13)
for any $\underline{x}\in S_{r,\underline{t}}$.
Fix a choice of $\underline{e}\in\mathcal{E}^{1}(Y)$ and set
$\mu_{\underline{e},\underline{t}}=\min\left\{\lambda_{\underline{e},\underline{t}},\left\lfloor\frac{Y-\log\underline{B}^{\underline{e}}}{S-\log\underline{B}^{\underline{t}}}\right\rfloor\right\}.$ (2.14)
Let $\underline{e}^{\prime}=\underline{e}-\lambda_{\underline{e},\underline{t}}\underline{t}$, where $\lambda_{\underline{e},\underline{t}}\in\mathbb{Z}_{\geqslant 0}$ is defined before the statement of the lemma. Notice that $\lambda_{\underline{e},\underline{t}}-\mu_{\underline{e},\underline{t}}\geqslant 0$ and that any monomial in $\underline{x}^{\underline{e^{\prime}}+(\lambda_{\underline{e},\underline{t}}-\mu_{\underline{e},\underline{t}})\underline{t}}g_{q}(\underline{x})^{\mu_{\underline{e},\underline{t}}}$ must have exponent vector belonging to $\mathcal{E}^{1}(Y)$, since the exponent of $x_{1}$ is $e_{1}^{\prime}=e_{1}<m_{1}$ and
$\log\underline{B}^{\underline{e^{\prime}}+(\lambda_{\underline{e},\underline{t}}-\mu_{\underline{e},\underline{t}})\underline{t}}+\mu_{\underline{e},\underline{t}}S=\log\underline{B}^{\underline{e},\underline{t}}+\mu_{\underline{e},\underline{t}}(S-\log\underline{B}^{\underline{t}})\leqslant Y.$
Moreover, we have $\underline{x}^{\underline{e^{\prime}}+(\lambda_{\underline{e},\underline{t}}-\mu_{\underline{e},\underline{t}})\underline{t}}g_{q}(\underline{x})^{\mu_{\underline{e},\underline{t}}}\equiv 0\bmod{q^{\mu_{\underline{e},\underline{t}}}}$ for all $\underline{x}\in S_{r,\underline{t}}$, by (2.13).
We proceed by performing elementary column operations to the column corresponding to $\underline{e}$, which will leave the value of $\Delta$ unchanged. On adding every other monomial of $\underline{x}^{\underline{e^{\prime}}+(\lambda_{\underline{e},\underline{t}}-\mu_{\underline{e},\underline{t}})\underline{t}}g_{q}(\underline{x})^{\mu_{\underline{e},\underline{t}}}$, we may assume that the column takes the form $\underline{x}^{\underline{e^{\prime}}+(\lambda_{\underline{e},\underline{t}}-\mu_{\underline{e},\underline{t}})\underline{t}}g_{q}(\underline{x})^{\mu_{\underline{e},\underline{t}}}$. But then each term of the column is divisible by $q^{\mu_{\underline{e},\underline{t}}}$. We repeat this argument for every $\underline{e}\in\mathcal{E}^{1}(Y)$, noticing that after every step, each column of the initial matrix can be written as a linear combination of the columns of the new matrix. Thus the fact that we already performed the procedure to other columns will not affect the procedure for any particular column. The statement of the lemma easily follows. ∎
Our next task is to get a useful bound for $\lambda$ in some particular cases.
Lemma 2.5.
Let $\underline{t}\in\{(0,0,0),(0,t_{2},0),(0,0,t_{3})\}$ and let $\lambda$ be given by (2.12). If $\underline{t}\neq(0,0,0)$, furthermore, assume that $S=\log\underline{B}^{\underline{t}}$. Then
$\lambda=\frac{2\sqrt{2}RE^{3/2}}{3}\left(1+O\left(\sqrt{\frac{\log B}{Y}}\right)\right),$
where $R$ is given by (2.10).
Proof.
We begin by assuming that $\underline{t}=(0,0,0)$, so that $\lambda_{\underline{e},\underline{t}}=\infty$ and $\log\underline{B}^{\underline{t}}=0$. We then have
$\lambda=\sum_{\underline{e}\in\mathcal{E}^{1}(Y)}\left\lfloor\frac{Y-\log\underline{B}^{\underline{e}}}{S}\right\rfloor$ $=\sum_{\underline{e}\in\mathcal{E}^{1}(Y)}\left(\frac{Y-\log\underline{B}^{\underline{e}}}{S}+O(1)\right)$
$=\frac{Y}{S}\#\mathcal{E}^{1}(Y)-\frac{1}{S}\sum_{\underline{e}\in\mathcal{E}^{1}(Y)}\log\underline{B}^{\underline{e}}+O\left(\#\mathcal{E}^{1}(Y)\right).$
Moreover,
$\sum_{e_{2}\log B_{2}+e_{3}\log B_{3}\leqslant Y_{1}}1=\frac{Y_{1}^{2}}{2\log B_{2}\log B_{3}}\left(1+O\left(\frac{\log B}{Y_{1}}\right)\right),$
by [8, Eq. (5.6)]. Thus we obtain
#$\#\mathcal{E}^{1}(Y)$ $=\sum_{e_{1}<m_{1}}\sum_{e_{2}\log B_{2}+e_{3}\log B_{3}\leqslant Y-e_{1}\log B_{1}}1$
$=\sum_{e_{1}<m_{1}}\frac{(Y-e_{1}\log B_{1})^{2}}{2\log B_{2}\log B_{3}}\left(1+O\left(\frac{\log B}{(Y-e_{1}\log B_{1})}\right)\right)$
$=\frac{m_{1}Y^{2}}{2\log B_{2}\log B_{3}}\left(1+O\left(\frac{\log B}{Y}\right)\right),$
since $e_{1}\log B_{1}<m_{1}\log B_{1}\ll\log B\ll Y$.
Next, we note that
$\sum_{e_{2}\log B_{2}+e_{3}\log B_{3}\leqslant Y_{1}}e_{2}$ $=\sum_{e_{2}\leqslant\frac{Y_{1}}{\log B_{2}}}\sum_{e_{3}\leqslant\frac{Y_{1}-e_{2}\log B_{2}}{\log B_{3}}}e_{2}$
$=\sum_{e_{2}\leqslant\frac{Y_{1}}{\log B_{2}}}\left\lfloor\frac{Y_{1}-e_{2}\log B_{2}}{\log B_{3}}+1\right\rfloor e_{2}$
$=\sum_{e_{2}\leqslant\frac{Y_{1}}{\log B_{2}}}\left(\frac{Y_{1}-e_{2}\log B_{2}}{\log B_{3}}\right)e_{2}+O\left(\frac{Y_{1}^{2}}{(\log B_{2})^{2}}\right)$
$=\frac{Y_{1}^{3}}{6(\log B_{2})^{2}\log B_{3}}+O\left(\frac{Y_{1}^{2}}{(\log B_{2})^{2}}\right),$
since
$\sum_{n\leqslant X}(X-n)n=\sum_{n\leqslant X}Xn-n^{2}=\frac{X^{3}}{2}-\frac{X^{3}}{3}+O(X^{2})=\frac{X^{3}}{6}+O(X^{2}),$
for any $X>0$. But then it follows that
$\sum_{\underline{e}\in\mathcal{E}^{1}(Y)}e_{2}=~{}$ $\sum_{e_{1}<m_{1}}\sum_{\begin{subarray}{c}e_{2}\log B_{2}+e_{3}\log B_{3}\leqslant Y-e_{1}\log B_{1}\end{subarray}}e_{2}$
[…]
Similarly,
$\sum_{\underline{e}\in\mathcal{E}^{1}(Y)}e_{3}=\frac{m_{1}Y^{3}}{6\log B_{2}(\log B_{3})^{2}}\left(1+O\left(\frac{\log B}{Y}\right)\right).$
Hence we deduce that
$\sum_{\underline{e}\in\mathcal{E}^{1}(Y)}\log\underline{B}^{\underline{e}}=~{}$ $\sum_{e_{1}<m_{1}}\sum_{\begin{subarray}{c}e_{2}\log B_{2}+e_{3}\log B_{3}\leqslant Y-e_{1}\log B_{1}\end{subarray}}e_{2}\log B_{2}+e_{3}\log B_{3}$
$+O\left(\#\mathcal{E}^{1}(Y)\log B_{1}\right)$
[…]
Returning to our initial expression for $\lambda$, and applying (2.9), we may now insert these estimates to conclude that
$\lambda$ $=\left(\frac{m_{1}Y^{3}}{2S\log B_{2}\log B_{3}}-\frac{m_{1}Y^{3}}{3S\log B_{2}\log B_{3}}\right)\left(1+O\left(\frac{\log B}{Y}\right)\right)$
$=\frac{m_{1}Y^{3}}{6S\log B_{2}\log B_{3}}\left(1+O\left(\frac{\log B}{Y}\right)\right)$
$=\frac{2\sqrt{2}m_{1}E^{3/2}(\log B_{1})^{3/2}(\log B_{2})^{1/2}(\log B_{3})^{1/2}}{6S(\log T_{\underline{m}})^{3/2}}\left(1+O\left(\sqrt{\frac{\log B}{Y}}\right)\right).$
This concludes the proof of the lemma in the case $\underline{t}=(0,0,0)$.
We proceed to discuss the case $\underline{t}=(0,t_{2},0)$, with $t_{2}\neq 0$, the case $\underline{t}=(0,0,t_{3})$ being identical. In this case we have $S-\log\underline{B}^{\underline{t}}=0$ in (2.12) and $\lambda_{\underline{e},\underline{t}}=\lfloor e_{2}/t_{2}\rfloor=e_{2}/t_{2}+O(1)$, in the notation before Lemma 2.4. But then it follows from (2) that
$\lambda=\sum_{\underline{e}\in\mathcal{E}^{1}(Y)}\lambda_{\underline{e},\underline{t}}$ $=\frac{1}{t_{2}}\sum_{\underline{e}\in\mathcal{E}^{1}(Y)}e_{2}+O(\#\mathcal{E}^{1}(Y))$
$=\frac{m_{1}Y^{3}}{6t_{2}(\log B_{2})^{2}\log B_{3}}\left(1+O\left(\frac{\log B}{Y}\right)\right).$
Recalling that $S=t_{2}\log B_{2}$, we may therefore conclude the statement of the lemma by applying (2.9), as previously. ∎
We are now ready to bring together our various results about the $p$-adic valuation of $\Delta$ in order to record a lower bound for it.
Lemma 2.6.
Let $\varepsilon>0$ and assume that $K_{\varepsilon}\geqslant 1$, in the notation of (2.1). Assume that $\gcd(q,g(0,0))=1$ and $\log r\pi_{X}\ll\log B$, and furthermore, that $B\gg_{\varepsilon}1$ and $Y\gg_{\varepsilon}\log B$ for sufficiently large implied constants that depend only on $\varepsilon$. Then either $\Delta=0$ or
$\log\Delta\geqslant\frac{2\sqrt{2}E^{3/2}}{3}\left(R\log q+\log\sqrt{E}r-\frac{\varepsilon}{2}\log B\right).$
Proof.
Since $\gcd(q,g(0,0))=1$, we may deduce from Lemmas 2.4 and 2.5 that $q^{\lambda}\mid\Delta$, where
$\lambda=\frac{2\sqrt{2}RE^{3/2}}{3}\left(1+O\left(\sqrt{\frac{\log B}{Y}}\right)\right).$
Combining this with Lemmas 2.1 and 2.2, we therefore deduce that
$\log\Delta$ $\geqslant\lambda\log q+\nu\log r+\sum_{\begin{subarray}{c}p\ll E^{1/2}\\ p\nmid qr\pi_{X}\end{subarray}}N_{p}\log p.$
Applying Lemma 2.3, we conclude that there exists $E_{1}\geqslant 0$ such that
$\log\Delta\geqslant\frac{2\sqrt{2}E^{3/2}}{3}\left(R\log q+\log\sqrt{E}r\right)-E_{1},$
with
$E_{1}\ll E^{3/2}\left(\sqrt{\frac{\log T_{\underline{m}}}{Y}}\log r+\log\log 3qr\pi_{X}+\sqrt{\frac{\log B}{Y}}R\log q\right).$
By (2.11), we have
$R\log q=\left(\frac{\prod_{i}\log B_{i}}{\log T_{\underline{m}}}\right)^{\frac{1}{2}}+\varepsilon\log B-\log K_{\varepsilon}=O_{\varepsilon}(\log B),$
since $K_{\varepsilon}\geqslant 1$. Moreover, $\log\log 3q=O(\log\log B)$. Hence
$E_{1}$ $\ll_{\varepsilon}E^{3/2}\left(\sqrt{\frac{\log B}{Y}}\log Br+\log\log 3qr\pi_{X}\right)$
$\ll_{\varepsilon}E^{3/2}\left(\log\log B+\log B\sqrt{\frac{\log B}{Y}}\right),$
since we have assumed that $\log r\pi_{X}\ll\log B$. On assuming that $B$ and $Y/\log B$ are bigger than a sufficiently large constant depending only on $\varepsilon$, we readily conclude that $E_{1}<\frac{\varepsilon}{2}E^{3/2}\log B$, from which the statement of the lemma is clear. ∎
Lemma 2.6 is a critical ingredient in Theorem 1.1. For Theorem 1.2 we will need the following variant.
Lemma 2.7.
Let $\varepsilon>0$ and assume that $K_{\varepsilon}\geqslant 1$, in the notation of (2.1). Assume that $B_{1}=B_{2}=B_{3}=B$ and that $q$ satisfies the hypothesis (1.5). Assume that $\log r\pi_{X}\ll\log B$, and furthermore, that $B\gg_{\varepsilon}1$ and $Y\gg_{\varepsilon}\log B$ for sufficiently large implied constants that depend only on $\varepsilon$. Then either $\Delta=0$ or
$\log\Delta\geqslant\frac{2\sqrt{2}E^{3/2}}{3}\left(R\log q+\log\sqrt{E}r-\frac{\varepsilon}{2}\log B\right).$
Proof.
Let us suppose that $l=\deg g$ and let $p^{j}\|q$. If $p\nmid g(0,0)$ then we deduce from Lemmas 2.4 and 2.5 that $p^{j\lambda}\mid\Delta$, where
$\lambda=\frac{2\sqrt{2}RE^{3/2}}{3}\left(1+O\left(\sqrt{\frac{\log B}{Y}}\right)\right).$
If $p\mid g(0,0)$, on the other hand, then (1.5) implies that $p\nmid g_{0}(1,0)$ or $p\nmid g_{0}(0,1)$ and so we can apply Lemma 2.4 with $\underline{t}\in\{(0,l,0),(0,0,l)\}$. But then Lemma 2.5 leads to precisely the same conclusion, since $S=l\log B$. Thus we may conclude that $q^{\lambda}\mid\Delta$ and the remainder of the argument proceeds exactly as in the proof of Lemma 2.6. ∎
Building on the previous two results we can prove the following result, in which we recall the notation (2.6) for $S_{r,\underline{t}}$.
Lemma 2.8.
Let $\varepsilon>0$ and assume that $K_{\varepsilon}\geqslant 1$, in the notation of (2.1). Assume that $\gcd(q,g(0,0))=1$, or that $B_{1}=B_{2}=B_{3}=B$ and $q$ satisfies (1.5). Assume that $\log\pi_{X}\ll\log B$ and $B\gg_{\varepsilon}1$, for a sufficiently large implied constant depending only on $\varepsilon$. Let $r_{1}<\dots<r_{u}$ be a sequence of primes that are coprime with $q$ and such that the reduction modulo $r_{i}$ of $X$ is geometrically irreducible. Let $r$ be given by (2.5) and assume that $\log r\ll\log B$. For each $1\leqslant i\leqslant u$, let $\underline{t}_{i}\in X(\mathbb{F}_{r_{i}})$ be a non-singular point of the reduction modulo $r_{i}$ of $X$. Then there exists a polynomial $f_{\underline{t}}\in\mathbb{Z}[\underline{x}]$, which is coprime with $f$, such that the points in $S_{r,\underline{t}}$ are all zeros of $f_{\underline{t}}$. Moreover, the polynomial has degree $O_{\varepsilon}(\max\{1,K_{\varepsilon}/r\}\log B/\log B^{\prime})$.
Proof.
Take $Y$ to be the smallest positive number such that $Y/\log B$ is bigger than a constant depending only on $\varepsilon$, as in Lemmas 2.6 or 2.7, and such that $\sqrt{E}r>K_{\varepsilon}$. Let $\Delta$ be the determinant of an $E\times E$ minor of the matrix $\mathbf{M}$ as defined in (2.7). If $\Delta=0$ for all $E\times E$ minors, then $\operatorname{rank}\mathbf{M}<E$ and there is a $\mathbb{Z}$-linear combination $f_{\underline{t}}(\underline{x})=\sum_{\underline{e}\in\mathcal{E}(Y)}a_{\underline{e}}\underline{x}^{\underline{e}}$ of monomials in $\mathcal{E}(Y)$, such that the points in $S_{r,\underline{t}}$ all vanish on $f_{\underline{t}}$. Moreover, as we proved at the beginning of the section, $f_{\underline{t}}$ will be coprime to $f$. Our choice of $Y$ ensures that
$1<\frac{\sqrt{E}r}{K_{\varepsilon}}=\sqrt{\frac{\log T_{\underline{m}}}{\log\prod_{i}B_{i}}}\cdot\frac{Y}{\sqrt{2}}\left(1+O\left(\sqrt{\frac{\log T_{\underline{m}}}{Y}}\right)\right)\frac{r}{K_{\varepsilon}},$
by (2.3). Since $Y$ is defined to be the smallest positive integer satisfying the constraints imposed in its definition, so it follows that $Y=O_{\varepsilon}(\max\{1,K_{\varepsilon}/r\}\log B)$. Finally, the degree of $f_{\underline{t}}$ is bounded by $\lfloor Y/\log B^{\prime}\rfloor$, whence
$\deg f_{\underline{t}}\ll Y/\log B^{\prime}\ll_{\varepsilon}\max\{1,K_{\varepsilon}/r\}\log B/\log B^{\prime},$
as claimed.
It remains to consider the possibility that $\Delta\neq 0$ for an $E\times E$ minor. But then it follows from combining (2.11) with Lemmas 2.6 or 2.7 that
$\log\Delta$ $\geqslant\frac{2\sqrt{2}E^{\frac{3}{2}}}{3}\left(R\log q+\log\sqrt{E}r-\frac{\varepsilon}{2}\log B\right)$
$=\frac{2\sqrt{2}E^{\frac{3}{2}}}{3}\left(-\log K_{\varepsilon}+\log\sqrt{E}r+\sqrt{\frac{\prod_{i}\log B_{i}}{\log T_{\underline{m}}}}+\frac{\varepsilon}{2}\log B\right).$
However, (2.8) and (2.9) combine to yield
$\log\Delta\leqslant\frac{2\sqrt{2}E^{\frac{3}{2}}}{3}\sqrt{\frac{\prod_{i}\log B_{i}}{\log T_{\underline{m}}}}\left(1+O\left(\sqrt{\frac{\log B}{Y}}\right)\right).$ (2.16)
Hence
$\log\Delta\leqslant\frac{2\sqrt{2}E^{\frac{3}{2}}}{3}\left(\sqrt{\frac{\prod_{i}\log B_{i}}{\log T_{\underline{m}}}}+\frac{\varepsilon}{2}\log B\right),$
which contradicts the assumption $\sqrt{E}r>K_{\varepsilon}$. ∎
We now have everything in place to complete the proof of Theorems 1.1 and 1.2. Let $\varepsilon>0$ and recall the notation $B^{\prime}$ and $B$ from (1.2). Clearly, we may work under the assumption that $B\gg_{\varepsilon}1$, for a sufficiently large implied constant that depends on $\varepsilon$.
We begin by dealing with the case $K_{\varepsilon}\leqslant 1$, in the notation of (2.1). (In particular, it will then follow that $KB^{\varepsilon}\leqslant 1$ in (1.3) and (1.4).) In this case we fix $r=1$ and we choose $Y\gg_{\varepsilon}\log B$ such that (2.4) holds. Since $K_{\varepsilon}\leqslant 1$, we deduce from (2.11) that
$R\log q\geqslant\sqrt{\frac{\prod_{i}\log B_{i}}{\log T_{\underline{m}}}}+\varepsilon\log B.$
Let $\Delta$ be the determinant of an $E\times E$ minor of the matrix $\mathbf{M}$ defined in (2.7). Here $r=1$, so that $\underline{t}$ is the empty vector and
$S_{1,\underline{t}}=\left\{\underline{x}\in X(\mathbb{Z}):|x_{i}|\leqslant B_{i},~{}g(\underline{x})\equiv 0\bmod q\right\}.$
The first few lines of the proofs of Lemmas 2.6 and 2.7 ensure that $q^{\lambda}\mid\Delta$, under the assumptions of Theorems 1.1 and 1.2, with
$\lambda=\frac{2\sqrt{2}RE^{\frac{3}{2}}}{3}\left(1+O\left(\sqrt{\frac{\log B}{Y}}\right)\right).$
Hence
$\log\Delta$ $\geqslant\frac{2\sqrt{2}RE^{\frac{3}{2}}}{3}\left(1+O\left(\sqrt{\frac{\log B}{Y}}\right)\right)\log q$
$\geqslant\frac{2\sqrt{2}E^{\frac{3}{2}}}{3}\left(\sqrt{\frac{\prod_{i}\log B_{i}}{\log T_{\underline{m}}}}+\varepsilon\log B\right)\left(1+O\left(\sqrt{\frac{\log B}{Y}}\right)\right),$
if $\Delta\neq 0$. This inequality is incompatible with the competing upper bound (2.16), if $Y\gg_{\varepsilon}\log B$ for a sufficiently large implied constant. Hence $\Delta=0$ and we can construct an auxiliary polynomial $f_{1}(\underline{x})=\sum_{\underline{e}\in\mathcal{E}(Y)}a_{\underline{e}}\underline{x}^{\underline{e}}$, as in the proof of Lemma 2.8, which has all the properties required of it in the case $K_{\varepsilon}\leqslant 1$.
We are now ready to deal with the case $K_{\varepsilon}\geqslant 1$. If $\log\pi_{X}\ll\log B$ fails to hold, then the desired conclusion follows from [14, Lemma 1.9]. Thus we may proceed under the assumption that $\log\pi_{X}\ll\log B$. But then every assumption of Lemma 2.8 holds and the proof is now identical to the one provided by Salberger [14, Lemma 3.2]. This completes the statement of Theorem 1.1 and 1.2.
Remark 2.9.
It is possible to prove a version of Theorem 1.1 in which the assumption $\gcd(q,g(0,0))=1$ is replaced with the weaker assumption that $q$ is coprime to the content of $g$. However, this would lead to a more complicated expression for $K$ in (2.1) and will not be carried out here.
3. Counting points on quadric surfaces
In this section we prove Corollary 1.3, which concerns an upper bound for the counting function $N_{1}(B)$ defined in (1.6). Let $S\subset\mathbb{A}^{3}$ be the quadric surface
$a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}=n$
for $a_{1},a_{2},a_{3},n\in\mathbb{Z}$ such that $-a_{1}a_{2}a_{3}n\neq\square$ and $\gcd(a_{1},a_{2},a_{3},n)=1$. The condition $-a_{1}a_{2}a_{3}n\neq\square$ precisely ensures that $S$ contains no lines defined over $\mathbb{Q}$. Indeed, if $X\subset\mathbb{P}^{3}$ is the compactification of $S$, then $\operatorname{Pic}(X)\cong\mathbb{Z}^{2}$ if and only if $X(\mathbb{Q})\neq\emptyset$ and the underlying quadratic form has square determinant. Thus any line contained in $S$ contains only $O(1)$ points.
When $\max\{|n|,|\underline{a}|\}\gg B^{20}$, for a sufficiently large implied constant, it follows from [5, Lemma 5] (which is based on an idea of Heath-Brown [7, Thm. 4]), that there exists an auxiliary quadratic polynomial $q\in\mathbb{Z}[x_{1},x_{2},x_{3}]$, which is coprime to the polynomial defining $S$, and such that the points counted by $N_{1}(B)$ are also zeros of $q$. The intersection of $S$ with the quadric $q=0$ produces at most $4$ irreducible curves and Bombieri–Pila [1, Thm. 5] implies that each irreducible curve makes a contribution $O_{\varepsilon}(B^{1/2+\varepsilon})$, since any lines have only $O(1)$ points. In this way it follows that
$N_{1}(B)\ll_{\varepsilon}B^{\frac{1}{2}+\varepsilon}$
if $\max\{|n|,|\underline{a}|\}\gg B^{20}$, which is satisfactory. We may therefore proceed under the assumption that $\max\{|n|,|\underline{a}|\}\ll B^{20}$.
Suppose without loss of generality that $|a_{1}|=|\underline{a}|$. We shall rework Section 2 with $q=|a_{1}|$ and $f(\underline{x})=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}-n$ and $g(\underline{x})=a_{2}x_{2}^{2}+a_{3}x_{3}^{2}-n$. Notice that $g(\underline{x})\equiv 0\bmod{q}$ for any $\underline{x}\in S(\mathbb{Z})$. Moreover, we have
$\gcd(q,g(0,0),g_{0}(1,0),g_{0}(0,1))=\gcd(a_{1},n,a_{2},a_{3})=1.$
Let $\pi_{S}=\left\lvert 2a_{1}a_{2}a_{3}n\right\rvert$ and note that the reduction modulo $p$ of $S$ is non-singular and geometrically integral for any $p\nmid\pi_{S}$. Moreover, we have $\log\pi_{S}\ll\log B$. We shall take $B_{1}=B_{2}=B_{3}=B$ and $\underline{m}=(2,0,0)$, noting that $(2,0,0)$ is the maximum among the exponents appearing in $f$ with non-zero coefficient for the lexicographic order.
The following result is extracted from work of Salberger [13] and will act as a proxy for Lemma 2.1 in the present setting.
Lemma 3.1.
There exists $\nu\in\mathbb{N}$ such that $r^{\nu}\mid\Delta$, with
$\nu=E^{\frac{3}{2}}\left(1+O\left(E^{-1/2}\right)\right).$
Proof.
When $B_{1}=B_{2}=B_{3}=B$, it is natural to take $Y=n\log B$ in the definition of $\mathcal{E}(Y)$, where $n$ is the maximum allowable degree of the monomials indexed by the set. In particular, we have $E=\#\mathcal{E}(Y)=Q(n)$, where
$Q(n)=\binom{n+3}{3}-\binom{n+1}{3}=n^{2}+O(n).$
On taking $N=3$ in the remark after Lemma 16.11 in [13], we deduce that $r^{\nu}\mid\Delta$, with $\nu=n^{3}+O(n^{2})$. The statement of the lemma easily follows. ∎
The next result is a refinement of Lemma 2.2.
Lemma 3.2.
Let $p\nmid\pi_{S}$. There exists $N_{p}\in\mathbb{N}$ such that $p^{N_{p}}\mid\Delta$, with
$N_{p}\geqslant\frac{E^{3/2}}{p}+O\left(E+E^{3/2}p^{-3/2}\right).$
Proof.
The proof is identical to the one provided by [14, Lemmas 1.4 and 1.5], but modified to account for the new lower bound for the exponent that follows from the remark after [13, Lemma 16.11], which is valid since the reduction of $S$ modulo $p$ is non-singular for any $p\nmid\pi_{S}$. ∎
In what follows we take
$K^{\prime}=q^{-\frac{1}{6}}B^{\frac{2}{3}},$ (3.1)
recalling that $q=|\underline{a}|$. Let $K_{\varepsilon}^{\prime}=K^{\prime}B^{\varepsilon}$. We are now ready to prove our analogue of Lemma 2.8 using the new bounds.
Lemma 3.3.
Let $\varepsilon>0$ and assume that $K^{\prime}_{\varepsilon}\geqslant 1$. Assume that $B\gg_{\varepsilon}1$, for a sufficiently large implied constant depending only on $\varepsilon$. Let $r_{1}<\dots<r_{u}$ be a sequence of primes that are coprime with $q$ and such that the reduction modulo $r_{i}$ of $S$ is geometrically irreducible. Let $r$ be given by (2.5) and assume that $\log r\ll\log B$. For each $1\leqslant i\leqslant u$, let $\underline{t}_{i}\in S(\mathbb{F}_{r_{i}})$ be a non-singular point of the reduction modulo $r_{i}$ of $S$. Then there exists a polynomial $f_{\underline{t}}\in\mathbb{Z}[\underline{x}]$, which is coprime with $f$, such that the points in $S_{r,\underline{t}}$ are all zeros of $f_{\underline{t}}$. Moreover, the polynomial has degree $O_{\varepsilon}(\max\{1,K^{\prime}_{\varepsilon}/r\})$.
Proof.
The proof is similar to Lemma 2.8, but using Lemmas 3.1 and 3.2 instead of Lemmas 2.1 and 2.2. We present the argument in full for the sake of clarity. Since $B_{1}=B_{2}=B_{3}=B$ we are working with $Y=n\log B$ for a large parameter $n$. We assume that $n\gg_{\varepsilon}1$ is the smallest integer such that $\sqrt{E}r>K_{\varepsilon}^{\prime}$, where $E=Q(n)=n^{2}+O(n)$, as in the proof of Lemma 3.1.
Let $\Delta$ be the determinant of an $E\times E$ minor of the matrix $\mathbf{M}$ as defined in (2.7). If $\Delta=0$ for all $E\times E$ minors, then $\operatorname{rank}\mathbf{M}<E$ and there is a $\mathbb{Z}$-linear combination $f_{\underline{t}}(\underline{x})=\sum_{\underline{e}\in\mathcal{E}(Y)}a_{\underline{e}}\underline{x}^{\underline{e}}$ of monomials in $\mathcal{E}(Y)$, such that the points in $S_{r,\underline{t}}$ all vanish on $f_{\underline{t}}$. Moreover, $f_{\underline{t}}$ will be coprime to $f$ and our choice of $Y$ ensures that
$1<\frac{\sqrt{E}r}{K_{\varepsilon}^{\prime}}=n\left(1+O\left(\frac{1}{n}\right)\right)\frac{r}{K_{\varepsilon}^{\prime}}.$
Since $n$ is defined to be the smallest positive integer satisfying the constraints imposed in its definition, so it follows that $\deg f_{\underline{t}}\leqslant n\ll_{\varepsilon}\max\{1,K_{\varepsilon}^{\prime}/r\},$ as claimed.
It remains to consider the possibility that $\Delta\neq 0$ for an $E\times E$ minor. In this case we wish to apply the variant of Lemma 2.7 tailored to the situation at hand. We claim that
$\log\Delta$ $\geqslant E^{3/2}\left(\frac{\log q}{6}+\log\sqrt{E}r-\frac{\varepsilon}{2}\log B\right).$
To see this, we first apply Lemma 2.5 to deduce that $q^{\lambda}\mid\Delta$, with
$\lambda=\frac{E^{3/2}}{6}\left(1+O\left(E^{-1/4}\right)\right),$
since $R=1/(4\sqrt{2})$ in (2.10) and $n\gg\sqrt{E}$. Combining this with Lemmas 3.1 and 3.2, and then with the modified version of Lemma 2.3, we therefore deduce that
$\log\Delta$ $\geqslant\lambda\log q+\nu\log r+\sum_{\begin{subarray}{c}p\ll E^{1/2}\\ p\nmid qr\pi_{X}\end{subarray}}N_{p}\log p$
$\geqslant E^{3/2}\left(\frac{\log q}{6}+\log\sqrt{E}r\right)-E_{1},$
for some $E_{1}$ satisfying
$0\leqslant E_{1}\ll_{\varepsilon}E^{3/2}\left(E^{-1/4}\log q+E^{-1/2}\log r+\log\log 3qr\pi_{S}\right).$
Clearly $E_{1}<\frac{\varepsilon}{2}E^{3/2}\log B$ if $n\gg_{\varepsilon}1$, since $\log q\ll\log B$ and we have assumed that $\log r\pi_{S}\ll\log B$. This establishes the claim. However, it follows from (2.16) that
$\log\Delta\leqslant\frac{2E^{\frac{3}{2}}}{3}\left(\log B+\frac{\varepsilon}{2}\log B\right),$
which contradicts the assumption $\sqrt{E}r>K_{\varepsilon}^{\prime}$. ∎
We are now ready to conclude the proof of Corollary 1.3. Recall the definition (3.1) of $K^{\prime}$, in which we recall that $q=|\underline{a}|$. Our aim is to establish the existence of polynomials $f_{1},\ldots,f_{J}\in\mathbb{Z}[x_{1},x_{2},x_{3}]$, and a finite collection of points $Z\subset S$, such that
  1. (1)
    $J=O_{\varepsilon}(\max\{1,K^{\prime}B^{\varepsilon}\})$;
  2. (2)
    each $f_{j}$ is coprime to $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}-n$ and has degree $O_{\varepsilon}(1)$, for $j\leqslant J$;
  3. (3)
    #$\#Z=O_{\varepsilon}(\max\{1,K^{\prime}B^{\varepsilon}\}^{2})$; and
  4. (4)
    for each $\underline{x}$ counted by $N_{1}(B)$, with $\underline{x}\not\in Z$, there exists $j\leqslant J$ such that $f_{j}(x_{1},x_{2},x_{3})=0.$
If $K^{\prime}B^{\varepsilon}\leqslant 1$ then the proof is identical to the corresponding case treated in the proof of Theorem 1.2. Alternatively, if $K^{\prime}B^{\varepsilon}\geqslant 1$ then the same proof goes through, making sure to use Lemma 3.3 instead of Lemma 2.8.
It remains to handle the points on the $J$ curves contained in $S$. By Bombieri–Pila [1, Thm. 5], each irreducible curve makes a contribution $O_{\varepsilon}(B^{1/2+\varepsilon})$, since the condition $-a_{1}a_{2}a_{3}n\neq\square$ ensures that any lines contain $O(1)$ points. Thus the overall contribution from the curves is $O_{\varepsilon}(\max\{1,K^{\prime}\}B^{1/2+2\varepsilon})$ and the statement of Corollary 1.3 follows on redefining $\varepsilon$.
4. Sums of unlike powers
In this section we prove Corollary 1.4. In fact we shall begin by proving a more general result.
4.1. Counting points on affine threefolds
Consider the threefold $X\subseteq\mathbb{A}^{4}$ defined by the polynomial
$f(x_{1},x_{2},x_{3},x_{4})=(\alpha x_{1}+\beta x_{4}+\gamma)h(x_{1},x_{2},x_{3},x_{4})+g(x_{2},x_{3}),$ (4.1)
where $\alpha,\beta,\gamma\in\mathbb{Z}$ and $g,h$ are polynomials with integer coefficients, with $\deg(g)=l$ and $\deg(h)=\deg_{x_{1}}(h)=k$. We further assume that $\beta\neq 0$. We let $X^{\circ}$ be the set of points in $X$ that do not belong to a rational line or a rational conic contained in $X$. We will use Theorem 1.2 to bound the quantity $N^{\circ}(B)=\#\{\underline{x}\in X^{\circ}(\mathbb{Z}):|\underline{x}|\leqslant B\},$ as follows.
Theorem 4.1.
Assume that the surface $X\cap\{\alpha x_{1}+\beta x_{4}+\gamma=q\}$ is geometrically irreducible for any non-zero $q\in\mathbb{Z}$. Assume that $\gcd(g(0,0),g_{0}(1,0),g_{0}(0,1))=1$, where $g_{0}$ is the top degree part of $g$. Then, for any $\varepsilon>0$, we have
$N^{\circ}(B)\ll_{\varepsilon}\max\{|\alpha|,|\beta|,|\gamma|\}B^{\varepsilon}\left(B^{\frac{4}{3}+\frac{1}{\sqrt{k}}(1-\frac{1}{2l})}+B^{1+\frac{2}{\sqrt{k}}(1-\frac{1}{2l})}\right).$
Before proving the theorem, we need a result from elementary number theory.
Lemma 4.2.
Given $-1<\alpha<0$ and $\varepsilon>0$, we have
$\sum_{u\leqslant X}(u/\gcd(u,n))^{\alpha}\ll_{\varepsilon}X^{\alpha+1}n^{\varepsilon}.$
Proof.
We have
$\sum_{u\leqslant X}(u/\gcd(u,n))^{\alpha}$ $=\sum_{d\mid n}\sum_{\begin{subarray}{c}u\leqslant X\\ \gcd(u,n)=d\end{subarray}}(u/d)^{\alpha}$
$\leqslant\sum_{d\mid n}\sum_{\begin{subarray}{c}u\leqslant X/d\end{subarray}}u^{\alpha}$
$\ll\sum_{d\mid n}\left(\frac{X}{d}\right)^{\alpha+1}$
$\ll n^{\varepsilon}X^{\alpha+1},$
by the divisor bound. ∎
Proof of Theorem 4.1.
Let $L=\max\{|\alpha|,|\beta|,|\gamma|\}$ and let $u\in\mathbb{Z}$. We write
$N_{u}^{\circ}(B)=\#\{\underline{x}\in X^{\circ}(\mathbb{Z}):|\underline{x}|\leqslant B,\alpha x_{1}+\beta x_{4}+\gamma=u\}.$
We have $N_{u}^{\circ}(B)=0$ unless $|u|\leqslant 3LB$, whence
$N^{\circ}(B)\leqslant\sum_{\left\lvert u\right\rvert\leqslant 3LB}N_{u}^{\circ}(B).$ (4.2)
The points in $X$ with $u=0$ are the points in $\{\alpha x_{1}+\beta x_{4}+\gamma=0\}\cap\{g(x_{2},x_{3})=0\}$. We can estimate the number of relevant $(x_{2},x_{3})$ using Bombieri–Pila [1, Thm. 5]. There are $O_{\varepsilon}(B^{1/3+\varepsilon})$ points that contribute, as we are not interested in points that lie on rational lines or conics. Since there are $O(B)$ choices of $(x_{1},x_{4})$ such that $\alpha x_{1}+\beta x_{4}+\gamma=0$, this therefore leads to the bound
$N_{0}^{\circ}(B)\ll_{\varepsilon}B^{4/3+\varepsilon}.$ (4.3)
Next assume that $u\neq 0$. It will be convenient to put $\deg_{x_{4}}(h)=k^{\prime}$ and to write
$h_{u}(x_{1},x_{2},x_{3})=\beta^{k^{\prime}}h(x_{1},x_{2},x_{3},(u-\alpha x_{1}-\gamma)/\beta).$
Since $\beta\neq 0$, we see that $h_{u}$ is a polynomial with integer coefficients and
$\beta^{k^{\prime}}h(x_{1},x_{2},x_{3},x_{4})=h_{u}(x_{1},x_{2},x_{3}),$
for any $\underline{x}$ counted by $N_{u}^{\circ}(B)$. Let $X_{u}\subseteq\mathbb{A}^{3}$ be the surface defined by the equation
$uh_{u}(x_{1},x_{2},x_{3})+\beta^{k^{\prime}}g(x_{2},x_{3})=0.$
This is geometrically irreducible by assumption. Moreover,
$N_{u}^{\circ}(B)\leqslant\#\{\underline{x}\in X_{u}^{\circ}(\mathbb{Z}):|\underline{x}|\leqslant B\},$
where $X_{u}^{\circ}$ is the set of points in $X_{u}$ that do not belong to any rational lines or conics in the surface. Put $q=|u|/\gcd(u,\beta^{k^{\prime}})$ and note that $g(x_{2},x_{3})\equiv 0\bmod u$ for any $(x_{1},x_{2},x_{3})\in X_{u}(\mathbb{Z})$. Moreover, the hypothesis (1.5) follows from the assumptions of the theorem. Thus
$N_{u}^{\circ}(B)\leqslant\#\{\underline{x}\in X_{u}^{\circ}(\mathbb{Z}):|\underline{x}|\leqslant B,~{}g(x_{2},x_{3})\equiv 0\bmod q\}$
and we can apply Theorem 1.2 to estimate this quantity.
We can take $\underline{m}=(k,0,0)$ since $\deg(h)=\deg_{x_{1}}(h)=k$. Indeed, $(k,0,0)$ is the maximum among the exponents appearing in the polynomial defining $X_{u}$ for the lexicographic order. Thus $K=B^{\frac{1}{\sqrt{k}}}q^{-\frac{1}{2\sqrt{k}l}}$ in (1.4). Then, for any $\varepsilon>0$, it follows that there is a set of $J=O_{\varepsilon}(\max\{1,KB^{\varepsilon}\})$ geometrically irreducible curves $C_{1},\dots,C_{J}\subset X_{u}$, with $\deg C_{j}=O_{\varepsilon}(1)$, such that there are $O_{\varepsilon}(\max\{1,KB^{\varepsilon}\}^{2})$ points counted by $N_{u}^{\circ}(B)$ that are not contained in any of these curves. When $\deg C_{j}\in\{1,2\}$, furthermore, we may assume that the relevant curve contains $O(1)$ integer points, since we are not interested in lines or conics defined over $\mathbb{Q}$. The overall contribution from the curves is therefore $O(JB^{1/3+\varepsilon})$, by [1, Thm. 5], which thereby shows that
$N_{u}^{\circ}(B)$ $\ll_{\varepsilon}\max\{1,KB^{\varepsilon}\}B^{\frac{1}{3}+\varepsilon}+\max\{1,KB^{\varepsilon}\}^{2}$
$\leqslant B^{\frac{1}{3}+\frac{1}{\sqrt{k}}+2\varepsilon}q^{-\frac{1}{2\sqrt{k}l}}+B^{\frac{2}{\sqrt{k}}+2\varepsilon}q^{-\frac{1}{\sqrt{k}l}}+B^{\frac{1}{3}+\varepsilon}.$
Combining this with (4.3) in (4.2), we conclude that
$N^{\circ}(B)$ $\ll_{\varepsilon}LB^{\frac{4}{3}+\varepsilon}+\sum_{\begin{subarray}{c}0<u\leqslant 3LB\end{subarray}}\left(\frac{B^{\frac{1}{3}+\frac{1}{\sqrt{k}}+2\varepsilon}}{(u/\gcd(u,\beta^{k^{\prime}}))^{\frac{1}{2\sqrt{k}l}}}+\frac{B^{\frac{2}{\sqrt{k}}+2\varepsilon}}{(u/\gcd(u,\beta^{k^{\prime}}))^{\frac{1}{\sqrt{k}l}}}\right).$
The statement of the theorem now follows from Lemma 4.2 and redefining the choice of $\varepsilon$. ∎
4.2. Proof of Corollary 1.4
We are now ready to prove Corollary 1.4. Let $N\in\mathbb{Z}$ be non-zero. Since $k$ is assumed to be odd, we can write
$x_{1}^{k}+x_{2}^{l}+x_{3}^{m}+x_{4}^{k}-N=(x_{1}+x_{4})h(x_{1},x_{4})+x_{2}^{l}+x_{3}^{m}-N,$
for a polynomial $h\in\mathbb{Z}[x_{1},x_{4}]$ of degree $k-1$. Let
$f(\underline{x})=(x_{1}+x_{4})h(x_{1},x_{4})+x_{2}^{l}+x_{3}^{m}-N.$
This is a polynomial of the form (4.1), with $g(x_{2},x_{3})=x_{2}^{l}+x_{3}^{m}-N$. In particular, $\gcd(g(0,0),g_{0}(1,0),g_{0}(0,1))=1$. We need to check that $X_{u}=X\cap\{x_{1}+x_{4}=u\}$ is geometrically irreducible for all non-zero $u\in\mathbb{Z}$. To do so, we pick $n\in\mathbb{Q}$ such that $M=N-uh(n,u-n)$ is non-zero. The polynomial $x_{2}^{l}+x_{3}^{m}-M$ is absolutely irreducible of degree $l>m\geqslant 2$, as one can see by applying the Eisenstein criterion over ¯$\overline{\mathbb{Q}}[x_{3}]$. Indeed, the polynomial $x_{3}^{m}-M\in\overline{\mathbb{Q}}[x_{3}]$ is square-free, since it has non-zero discriminant. Having checked all the hypotheses of Corollary 4.1, it now follows that the cardinality of points that do not lie on any lines or conics contained in $X$ is
$\ll_{\varepsilon}\left(B^{\frac{4}{3}+\frac{1}{\sqrt{k-1}}(1-\frac{1}{2l})}+B^{1+\frac{2}{\sqrt{k-1}}(1-\frac{1}{2l})}\right)B^{\varepsilon},$
since $l>m$. This is satisfactory for Corollary 1.4.
It remains to deal with the points that lie on lines and conics contained in the hypersurface $f=0$. For this we can adapt an argument of Newman and Slater [11].
Lemma 4.3.
Let $b\in\overline{\mathbb{Q}}^{*}$, let $l_{1},\dots,l_{r}\in\mathbb{N}$ and let $\gamma_{1}(t),\dots,\gamma_{r}(t)\in\overline{\mathbb{Q}}[t]$ such that
$\sum_{i=1}^{r}\gamma_{i}^{l_{i}}=b.$ (4.4)
Put $d_{i}=\deg(\gamma_{i})$ for $1\leqslant i\leqslant r$ and assume that $\gamma_{1}^{l_{1}},\dots,\gamma_{r}^{l_{r}}$ are linearly independent over $\overline{\mathbb{Q}}$. Then
$\max_{1\leqslant i\leqslant r}d_{i}l_{i}\leqslant(r-1)(d_{1}+\dots+d_{r})-\frac{r(r-1)}{2}.$
Proof.
Let $e_{i}=d_{i}l_{i}$ for $1\leqslant i\leqslant r$. Let $s_{i}=\max\{l_{i}-r+1,0\}$. Let $W$ be the Wronskian $W(\gamma_{1}^{l_{1}},\gamma_{2}^{l_{2}},\dots,\gamma_{r}^{l_{r}}).$ This is non-zero, since the polynomials are assumed to be linearly independent. On noticing that $\gamma_{1}^{s_{1}}\gamma_{2}^{s_{2}}\cdots\gamma_{r}^{s_{r}}$ must divide $W$, so it follows that
$\deg W\geqslant\sum_{1\leqslant i\leqslant r}s_{i}d_{i}.$
Let $e=\max_{1\leqslant i\leqslant r}e_{i}$ and assume without loss of generality that $e=e_{1}$. It follows from (4.4) that the Wronskian $W(b,\gamma_{2}^{l_{2}},\dots,\gamma_{r}^{l_{r}})$ is equal to $W$. Hence
$\deg W$ $\leqslant e_{2}+e_{3}+\dots+e_{r}-1-2-\dots-(r-1)$
$=\left(\sum_{1\leqslant i\leqslant r}e_{i}\right)-e-\frac{r(r-1)}{2}.$
But then
$e\leqslant\sum_{1\leqslant i\leqslant r}e_{i}-\sum_{1\leqslant i\leqslant r}s_{i}d_{i}-\frac{r(r-1)}{2}=\sum_{1\leqslant i\leqslant r}\left(d_{i}\left(l_{i}-s_{i}\right)-\frac{r-1}{2}\right).$
Clearly $l_{i}-s_{i}\leqslant l_{i}-(l_{i}-r+1)=r-1$, which leads to the statement of the lemma. ∎
Corollary 4.4.
Assume that $k\geqslant 13$ is odd, $k>l>m\geqslant 2$, and $N\in\overline{\mathbb{Q}}^{*}$. Then any lines or conics contained in the hypersurface $x_{1}^{k}+x_{2}^{l}+x_{3}^{m}+x_{4}^{k}=N$ must be contained in one of the subvarieties cut out by the union of equations
$V(x_{1}^{k}+x_{4}^{k})\cup V(x_{2}^{l}+x_{3}^{m})\cup V(x_{1}^{k}+x_{2}^{l}+x_{4}^{k})\cup V(x_{1}^{k}+x_{3}^{m}+x_{4}^{k}).$ (4.5)
Proof.
Assume that there is a line or conic contained in the hypersurface, so that
$\gamma_{1}(t)^{k}+\gamma_{2}(t)^{l}+\gamma_{3}(t)^{m}+\gamma_{4}(t)^{k}=N,$ (4.6)
with $\gamma_{i}\in\overline{\mathbb{Q}}[t]$ and $d_{i}=\deg(\gamma_{i})\leqslant 2$, for $1\leqslant i\leqslant 4$, which are not all constant. We begin by dispatching the case that one of the polynomials $\gamma_{1},\dots,\gamma_{4}$ is constant. We assume that $d_{4}=0$, the remaining cases being similar. If $d_{1}=2$, then the left hand side of (4.6) has degree $2k$. If $d_{1}=1$ and $d_{2}=2$, then the left hand side of (4.6) has degree $\max\{k,2l\}$. (Notice that $2l\neq k$, since $k$ is odd.) If $d_{1}=d_{2}=1$ and $d_{3}=2$, then the degree is $\max\{k,2m\}$. If $d_{1}=1$ and $d_{2},d_{3}\leqslant 1$, then the degree is $k$. In all of these cases the left hand side of (4.6) has strictly positive degree and so cannot be equal to the constant $N$. Hence $\gamma_{1}$ must also be constant. If $\gamma_{1}^{k}+\gamma_{4}^{k}=N$, then the curve lies on the surface cut out by the equation $x_{2}^{l}+x_{3}^{m}=0$ and we are done. If $\gamma_{1}^{k}+\gamma_{4}^{k}\neq N$, then $\gamma_{2}(t)^{l}+\gamma_{3}(t)^{m}=b$ for $b\in\overline{\mathbb{Q}}^{*}$. Moreover, $\gamma_{2}^{l}$ and $\gamma_{3}^{m}$ must be linearly independent, since at least one of $\gamma_{2}$ or $\gamma_{3}$ is non-constant. But then it follows from Lemma 4.3 that
$\max\{d_{2}l,d_{3}m\}\leqslant d_{2}+d_{3}-1.$
If $d_{2}=2$, then $2l\leqslant 1+d_{3}\leqslant 3$, which is impossible. If $d_{2}\leqslant 1$, then $d_{3}m\leqslant 1+d_{3}-1$, which is also impossible.
We may now proceed under the assumption that $d_{i}\in\{1,2\}$, for all $1\leqslant i\leqslant 4$. For notational convenience we write $(l_{1},l_{2},l_{3},l_{4})=(k,l,m,k)$. Let $r\in\mathbb{N}$ be the least integer such that we can write
$\gamma_{1}(t)^{l_{1}}+\gamma_{2}(t)^{l_{2}}+\gamma_{3}(t)^{l_{3}}+\gamma_{4}(t)^{l_{4}}=\sum_{j=1}^{r}a_{i_{j}}\gamma_{i_{j}}(t)^{l_{i_{j}}},$
for $\{i_{1},\dots,i_{r}\}\subseteq\{1,2,3,4\}$, with $a_{i_{j}}\in\overline{\mathbb{Q}}^{*}$ and $\gamma_{i_{1}}^{l_{i_{1}}},\dots,\gamma_{i_{r}}^{l_{i_{r}}}$ linearly independent.
If $r=4$ then Lemma 4.3 implies that
$\max\{d_{1},d_{4}\}k\leqslant\max_{1\leqslant j\leqslant 4}d_{j}l_{j}\leqslant 3(d_{1}+\dots+d_{4})-6.$
If $d_{1}=d_{4}=1$ then this implies that $k\leqslant 18-6=12$, which is a contradiction. If $\max\{d_{1},d_{4}\}=2$ then we obtain $182k\leqslant 24-6=$ 18, which is also impossible. If $r=3$ then at least one of $\gamma_{1}$ or $\gamma_{4}$ must belong to the sum. Assuming without loss of generality we fall in the former case, it follows from Lemma 4.3 that
$d_{1}k\leqslant\max_{1\leqslant j\leqslant 3}d_{i_{j}}l_{i_{j}}\leqslant 2(d_{1}+d_{i_{2}}+d_{i_{3}})-3.$
If $d_{1}=1$ this implies that $k\leqslant 10-3=7$, whereas if $d_{1}=2$ then we obtain $92k\leqslant 12-3=$ 9. Neither of these is possible. Suppose next that $r=2$. Then Lemma 4.3 implies that
$\max\{d_{i_{1}}l_{i_{1}},d_{i_{2}}l_{i_{2}}\}\leqslant d_{i_{1}}+d_{i_{2}}-1.$
But $\min\{l_{i_{1}},l_{i_{2}}\}\geqslant 2$, whence
$2\max\{d_{i_{1}},d_{i_{2}}\}\leqslant\max\{d_{i_{1}}l_{i_{1}},d_{i_{2}}l_{i_{2}}\}\leqslant d_{i_{1}}+d_{i_{2}}-1\leqslant 2\max\{d_{i_{1}},d_{i_{2}}\}-1,$
which is impossible. Finally, the case $r=1$ is impossible since we would then have $a_{i}\gamma_{i}(t)^{l_{i}}=N$, for a non-constant polynomial $\gamma_{i}$, for some $i\in\{1,\dots,4\}$. ∎
We may now conclude the proof of Corollary 1.4, for which it remains to control the contribution from lines or conics contained in the hypersurface $X\subset\mathbb{A}^{4}$ defined by
$x_{1}^{k}+x_{2}^{l}+x_{3}^{m}+x_{4}^{k}=N.$
Assuming that $k,l,m,N$ satisfy the hypotheses of the corollary, it follows from Corollary 4.4 that lines or conics on the hypersurface are all contained in the union of surfaces cut out by (4.5). Let $Y\subset\mathbb{A}^{3}$ be the surface obtained by intersecting $X$ with the first equation $x_{1}^{k}+x_{4}^{k}=0$ in (4.5). There are $O(B)$ choices with $|x_{1}|,|x_{4}|\leqslant B$ and $x_{1}^{k}+x_{4}^{k}=0$. We are then left with counting $|x_{2}|,|x_{3}|\leqslant B$ for which $x_{2}^{l}+x_{3}^{m}=N$. It follows from Eisenstein’s criterion over ¯$\overline{\mathbb{Q}}[x_{3}]$, as previously, that the underlying polynomial is absolutely irreducible of degree $l$. But then Bombieri–Pila [1, Thm. 5] reveals there to be $O_{\varepsilon}(B^{1/l+\varepsilon})$ choices for $x_{2},x_{3}$. In summary, the integral points $\underline{x}\in S(\mathbb{Z})$ make a satisfactory overall contribution $O_{\varepsilon}(B^{4/3+\varepsilon})$ to $N_{2}(B)$, since $l\geqslant 3$. The treatment of the other subvarieties defined by (4.5) is similar, which thereby completes the proof of Corollary 1.4.